In a trial of 325 patients who received​ 10-mg doses of a drug​ daily, 39 reported headache as a side effect. Obtain a point estimate for the population proportion of patients who received ​10-mg doses of a drug daily and reported headache as a side effect.

p=______ (Round to two decimal places as​ needed.)
Answer #1

x = 39

n = 325

We have to find the point estimate for the population proportion of patients who received ​10-mg doses of a drug daily and reported headache as a side effect.

p = x/n = 39/325 = 0.12